Find all numbers for which the rational expression is undefined.
2/6y+5
(Simplify your answer. Type an integer or a fraction. Use a comma to separate answers as needed. Type N is the answer is never undefined.)

Answer by stanbon(57361)

Find all numbers for which the rational expression is undefined.
2/6y+5
------------
When the denominator is zero, the rational expression is undefined.
6y+5 = 0
y = -5/6
(undefined when y = -5/6)
=============
Cheers,
Stan H.